Phillip’s bedroom is 15 feet by 16 feet with an 8-foot ceiling. There are two windows that are each 3 feet by
4 feet and two doors that are each 3 feet by 7 feet. How many square feet of wallpaper will Phillip need to fix
up the walls of his bedroom?

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)


Area of walls = 15*8 +16*8 =248 ft^2
less area of windows 2*(3*4)= 24 ft^2
Less area of doors 2*(3*7) = 42 ft ^2
248-66= 182 ft ^2
square feet of wallpaper = 182

Answer by ikleyn(53996)

.
Phillip’s bedroom is 15 feet by 16 feet with an 8-foot ceiling. There are two windows that are each 3 feet by
4 feet and two doors that are each 3 feet by 7 feet. How many square feet of wallpaper will Phillip need to fix
up the walls of his bedroom?
~~~~~~~~~~~~~~~~~~~~~~~~~~


        Calculations in the post by @mananth are incorrect.
        They are performed under the assumption that the bedroom has two walls - - - ha-ha-ha.
        I came to bring a correct solution.


Area of walls = 2*(15*8 + 16*8) = 496 ft^2

less area of windows 2*(3*4)= 24 ft^2

Less area of doors 2*(3*7) = 42 ft ^2

496 - 24 - 42 = 430 ft^2

square feet of wallpaper = 430

------------------

Solved correctly.